OpenStudy (anonymous):

lim e^-x + 5/4^1^x as x approaches infinity

OpenStudy (kirbykirby):

\[ \lim_{x\rightarrow -3} \frac{x^2}{x^2-9}\] If you try: \[ \lim_{x\rightarrow -3^-} \frac{x^2}{x^2-9}=\infty\] \[ \lim_{x\rightarrow -3^+} \frac{x^2}{x^2-9} =- \infty\] What does that say about the 2-sides limit ?

OpenStudy (mosaic):

If you put x = -3, the numerator is +9 but the denominator is 0. Therefore, the limit is infinity. But you have to see if the limit is +infinity or -infinity by approaching -3 from both sides. If the limits are different then the limit does not exist.

OpenStudy (anonymous):

it is approaching -3 from the right

OpenStudy (anonymous):

but i did not know how to put that in the question

OpenStudy (mosaic):

If you approach -3 from the left, you can start with x = -4 and approach -3. When x = -4, the numerator and denominator are both positive and so the limit when we approach x = -3 from the left is +infinity (which is what the first reply mentions in the middle section). When we approach -3 from the right, we can start with x = -2 which makes the numerator positive but denominator negative. Therefore, when we approach x = -3 from the right, the limit is -infinity. Since the left and right limits are different, the limit as x approaches -3 does not exist.

OpenStudy (anonymous):

thanks